5/x =0.2
Can you please slove x for me thank you !

Answer by waynest(281)

(5/x) = 0.2
x(5/x) = (0.2)x
5 = 0.2x
5/0.2 = 0.2x/0.2
25 = x
check:
(5/25) = 0.2
0.2 = 0.2